OpenStudy (anonymous):

Secants BE and CF intersect at point A inside the circle. Given that the measure of FAE = 60o and the measure of arc CE = 160 o, which is the measure of arc BF? 80 100 130 160

OpenStudy (anonymous):

Use this theorem to solve. If two secants intersect in the interior of a circle, then the measure of an angle formed is one-half the sum of the measures of the arcs intercepted by the angle and its vertical angle.

OpenStudy (anonymous):

i got it its 80 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i guessed

OpenStudy (anonymous):

OK, let's actually calculate it. Draw me a picture of a circle with two secants intersecting inside of the circle and then naming the angles.

OpenStudy (anonymous):

Are you drawing anything? I doesn't have to be pretty.

OpenStudy (anonymous):



OpenStudy (anonymous):

Angle CAE = 180-60 Which is 120 degrees. Angle BAF is a vertical angle of angle CAE So the sum of the arcs intercepted by angle CAE (160 deg) and vertical angle BAF 160 + M So angle CAE is equal to one-half of 160 + M. Which gets us 120=1/2(160 +M) Which gets us 240=160 + M So M = 80 degrees